# Masked Image Contrastive Learning for Efficient Visual Conceptual Pre-training

### v1

@@ 4.1 Scalability
To demonstrate the scalability of our MiCL for efficient conceptual pre-training, we access the efficiency and scaling of our model in Fig.3. It illustrates the pre-training hours related to various model sizes for different methods, with linear probing accuracy.
@@
Model Efficiency. MiCL presents highly scalable in contrast to prior methodologies, demanding reduced computational resources while delivering exemplary results, without relying on handcrafted data augmentations. Compared to reconstruction-based methods, such as MAE and I-JEPA, we necessitate fewer epochs and omit the need for context pixel-level reconstruction, thus resulting in significantly improved speed. On the flip side, concerning contrastive learning methods like MoCo v3 which rely on hand-crafted data augmentations for generating and processing multiple image views, MiCL dispenses with auxiliary modules like the momentum encoder akin to MoCo v3. The simple and straightforward framework of MiCL accelerates the pre-training. In particular, when the model becomes larger from ViT-B/16 to ViT-L/16, additional pre-training time is far away less than the MoCo v3.
Scaling model size. Moreover, our model leverages a scalable model size, resulting in more substantial performance enhancements with the larger model as illustrated in Fig.3. Compared to ViT-B/16, MiCL achieves nearly 4% improvement in linear probing with ViT-L/16, surpassing MoCo v3. It implies that we can efficiently train larger models to achieve better performance, within an acceptable timeframe.
@@ 4.2.1 Masked Ratio
Firstly, we conduct ablation experiments to discuss the impact of the overall masked ratio on the performance of conceptual pre-training, revealed in Table 2. To ensure clarity in presentation, we offer the visible ratio of each contrastive branch within a single forward pass of the mini-batch, which is a crucial factor for the cross-similarity within our MiCL framework. Likewise, the effective batch size is intertwined with the scalable masked ratios, with larger batch sizes showcasing performance enhancements for the model, as verified in Section 4.2.2. Thus, we attribute this improvement to the masked ratio.
@@
4.2.2 Large Batch Size for Contrastive Learning
Act. Bsz. Eff. Bsz. LIN FT
@@ 4.2.2 Large Batch Size for Contrastive Learning
512 2,048 77.9 85.8 533
512 4,096 77.7 85.7 533
450 1,800 77.4 85.7 559
256 1,024 74.4 84.9 586
128 512 71.6 84.6 613
Table 3: Ablation evaluation experiments on batch size. The results are based on the ImageNet-1K with the ViT-L/16. All methods are evaluated by pre-training hours, linear probing (LIN) and fine-tuning (FT). Different from other methods, MiCL is very dependent on the size of the mini-batch (denoted as Act. Bsz.), instead of effective batch size (denoted as Eff. Bsz.). So, we offer both parameters of batch size. Besides, pre-training hours on A100 are provided. The resolution of images is fixed to 224×224. Top-1 accuracy is used as the metric.
ViT models [8] are inherently computationally intensive, and training with large batches is a preferred strategy for handling large ViT models. Moreover, a sizable batch size is advantageous for achieving accuracy in contemporary self-supervised learning techniques. In particular, concerning contrastive learning methodologies that heavily lean on large batch sizes, ablation experiments are conducted to ascertain the influence of batch size on our masked image contrastive learning approach, as shown in Table 3. Diverging from other contrastive learning methods, our approach necessitates computing sample similarities within a mini-batch, making the actual batch size of the mini-batch more crucial than the effective batch size achievable through gradient accumulation. Thus, we provide both actual batch size and effective batch size in Table 3. The utilization of the ViT-L/16 model for validation reveals that a larger actual batch size correlates with improved model performance and efficiency, aligning with the consensus within the community [10, 30]. However, limited by computational resources, we are unaware of the maximum capacity of the actual batch size. Additionally, the model performance was not enhanced by the effective batch size achieved through gradient accumulation, while the actual batch size was more influential in this context.
@@ 4.2.3 MLP Head is Not You Need
Table 4: Ablation evaluation experiments on MLP head. The results are based on the ImageNet-1K with the ViT-L/16. All methods are evaluated by pre-training hours, linear probing (LIN) and fine-tuning (FT). Correspondingly, we give different effective batch sizes related to the MLP head within the pre-trained model. Besides, pre-training hours on A100 are provided. The resolution of images is fixed to 224×224. Top-1 accuracy is used as the metric.
Coupled with the ViT model, MLP head [22] is usually utilized to assist the pretext task in many contrastive learning methods [4, 6]. In this context, ViT is employed to learn the semantic features of the image, while the MLP head is utilized to manage the pretext classification task, thereby the assistance of the MLP head can enhance ViT’s learning capabilities as demonstrated in pioneer work. However, we argue that traditional contrastive learning methods require hand-crafted data augmentations for generating and processing multiple image views, which are based on the instance-level. While, MiCL endeavours to explore an alternative of using masked images to create diverse views with fine-grained conceptual differences on the token-level. Thus, we argue that MiCL does not require the MLP head to perform the instance-level classification tasks.
Subsequently, ablation experiments on the MLP head are conducted, opting for 2-layer and 3-layer MLPs, as well as no MLP head, as depicted in Table 4. Following MoCo v3 [6], the hidden layers of both 3-layer and 2-layer MLPs are 1024-d with GELU [13]. And the output layers of both MLPs are 512-d, without GELU. all layers in both MLPs have BN [14], following SimCLR [5]. Due to GPU memory limitations, we adjusted the batch size to 1,800 for our validation. It can be drawn from the fine-tuning results in Table 4, that MiCL can operate without an MLP head. Unlike conventional contrastive learning methods, it does not compromise the model’s performance. Conversely, the additional burden of extra auxiliary modules elevates both the cost and time of pre-training, with benefits that do not match the enhancements derived from increased batch sizes.
@@ 4.2.4 Concentration of Similarity Metric $\kappa$
Table 5: Ablation evaluation experiments on concentrate parameter of kappa. The results are based on the ImageNet-1K with the ViT-L/16. Fine-tuning (FT) results are provided. The resolution of images is fixed to 224×224. Top-1 accuracy is used as the metric.
Furthermore, following the T-SP [26], We perform ablation experiments to assess the influence of varying concentrations of the T-distributed adapter on the overall performance of pre-training model as shown in Table 5. This study involves five specific degrees of concentration, namely $\kappa=4$, $16$, $32$, $64$, and $128$. From the Table 5, we can infer that as kappa increases, the pre-training performance of the model improves progressively until reaching $\kappa=64$ with the fine-tuning result of 85.8 under ViT-L/16. We explain that, with the escalation of kappa, the model must extract more precise image semantic concepts when encountering positive samples, presenting a more challenging pretext task for model pre-training. However, an excessively large kappa will lead to the model converging too slowly, thereby impacting the effectiveness of the entire pre-training process.
@@ 4.3 Comparisons with previous results
ViT-B/16 800 - 83.2
BEiT [2] w/o ViT-L/16 800 - 85.2
ViT-B/16 1600 68.0 83.6
ViT-L/16 1600 76.0 85.9
MAE [12] w/o ViT-H/14 1600 77.2 86.9
ViT-B/16 1600 70.4 83.9
CAE [7] w/o ViT-L/16 1600 78.1 86.3
ViT-B/16 600 72.9 -
I-JEPA [1] w/o ViT-L/16 600 77.5 -
@@
SiameseIM [21] w/ ViT-B/16 1600 78.0 84.1
ccMIM [32] w/o ViT-B/16 800 68.9 84.2
ViT-B/16 800 - 85.3
ConMIM [29] w/ ViT-L/16 1600 - 86.5
ViT-B/16 1600 79.5 -
iBOT [33] w/ ViT-L/16 1200 81.0 84.0
Contrastive Learning
SimCLR v2 [5] w/ RN152 (2$\times$) 800 79.4 82.9
DINO [3] w/ ViT-B/16 1600 78.2 82.8
ViT-B/16 600 76.7 83.2
MoCo v3 [6] w/ ViT-L/16 600 77.6 84.1
ViT-B/16 800 74.2 83.4
MiCL w/o ViT-L/16 800 77.9 85.8
@@
To validate MiCL’s ability to acquire high-level conceptual representations without depending on hand-crafted data augmentations, we present comparison results of linear probing and fine-tuning classification under pre-training on ImageNet-1k. It is important to highlight that while masked images are utilized to diminish the semantic redundancy within images, our approach eschews the reconstruction of masked images in favor of leveraging contrastive loss to steer the entire network. Thus, our MiCL method is classified as contrastive learning.
@@
Table.6 exhibits the performance of our method under the fine-tuning and linear probing for ImageNet-1k classification. Our MiCL method has demonstrated outstanding performance in contrastive learning, notably obviating the need for intricate augmentations across multiple views. It validates the practicality and efficacy of masked images for generating diverse views encapsulating distinct fine-grained semantic concepts, which diminishes conceptual redundancy and expedites the conceptual pre-training process. Beyond that, it also stands out the significant competency of contrastive learning in extracting high-level semantic concepts.
@@
Compared to MIM methods, due to the absence of pixel-level image reconstruction in MiCL, the training process is completed in 800 epochs, in contrast to the 1600 epochs required for MAE [12], CAE [7]. It corroborates the contribution of our methodology from the perspective of efficiency and high-level semantic concept extraction. Moreover, our MiCL method is superior to BeiT [2] and I-JEPA [1] regarding the fine-tuning and linear probing results, also confirming the competency and robustness of our method.
@@
Concerning the fusion of MIM and CL, significant endeavours [21, 29, 33] are made to enhance pre-training performance on downstream tasks. However, they result in inefficiencies due to increased training iterations, augmented data, auxiliary modules, and diverse loss combinations. For instance, SiameseIM [21], ConMIM [29] and iBOT [33] leverage both hand-crafted data augmentation and more training epochs. Despite our model also integrating masking strategy with contrastive learning, it does not rely on hand-crafted view data augmentations and additional auxiliary modules, reconciling the divergence between efficient visual representation and effective conceptual pre-training. Besides, our model MiCL also achieves competitive results on downstream tasks of linear probing and fine-tuning, and more importantly, shows promising scaling behaviour.
More comparison results of downstream tasks are provided in the supplementary.
5 Conclusions
@@
In this paper, we present MiCL, a novel, simple and effective pre-training paradigm for visual conceptual representation. The masking strategy generates various views with nuanced semantic distinctions, subsequently leveraging contrastive learning to learn the agreement for classification within a mini-batch. In this way, we manage to avoid I) semantic conceptual redundancy within an image, II) reconstruction of images, III) hand-crafted data augmentations and IV) additional auxiliary modules, leading to improvement of efficiency and scalability. Experiments showcase the efficiency and scalability of our method, yielding competitive results when compared with previous approaches. Furthermore, the ablation experiments have sparked new considerations for future pre-training paradigms.

### v2

@@ 3.1 Scalability
To demonstrate the scalability of our OCL for efficient conceptual pre-training, we access the efficiency and scaling of our model in Fig.3. It illustrates the pre-training hours related to various model sizes for different methods, with linear probing accuracy.
@@
OCL is highly scalable compared to previous methods, requiring less computational resources while achieving strong results and without relying on handcrafted data augmentations. Unlike reconstruction-based approaches like MAE and I-JEPA, OCL needs fewer training epochs and avoids the need for pixel-level reconstruction, significantly improving training speed. In contrast to contrastive learning methods such as MoCo v3, which depend on handcrafted augmentations and complex architectures to generate and process multiple image views, OCL eliminates the need for auxiliary modules like momentum encoders. This simplicity makes OCL’s framework more efficient and accelerates pre-training. For example, when scaling up from ViT-B/16 to ViT-L/16, OCL requires far less additional pre-training time compared to MoCo v3.
Scaling model size. Moreover, our model leverages a scalable model size, resulting in more substantial performance enhancements with the larger model as illustrated in Fig.3. Compared to ViT-B/16, OCL achieves nearly 4% improvement in linear probing with ViT-L/16, surpassing MoCo v3. It implies that we can efficiently train larger models to achieve better performance, within an acceptable timeframe.
@@ 3.2.1 Masked Ratio
Firstly, we conduct ablation experiments to discuss the impact of the overall masked ratio on the performance of conceptual pre-training, revealed in Table 2. To ensure clarity in presentation, we offer the visible ratio of each contrastive branch within a single forward pass of the mini-batch, which is a crucial factor for the cross-similarity within our OCL framework. Likewise, the effective batch size is intertwined with the scalable masked ratios, with larger batch sizes showcasing performance enhancements for the model, as verified in Section 3.2.2. Thus, we attribute this improvement to the masked ratio.
@@
3.2.2 Large Batch Size for Effective Contrastive Representation
Effective
Batch Size
LIN FT
@@ 3.2.2 Large Batch Size for Effective Contrastive Representation
2,048 77.9 85.8 533
4,096 77.7 85.7 533
1,800 77.4 85.7 559
1,024 74.4 84.9 586
512 71.6 84.6 613
Table 3: Ablation evaluation experiments on batch size. The results are based on the ImageNet-1K with the ViT-L/16. All methods are evaluated by pre-training hours, linear probing (LIN) and fine-tuning (FT). Meanwhile, pre-training hours on A100 are provided. The resolution of images is fixed to 224×224. Top-1 accuracy is used as the metric.
ViT models [10] are inherently computationally intensive, and training with large batches is a preferred strategy for handling large ViT models. Moreover, a sizable batch size is advantageous for achieving accuracy in contemporary self-supervised learning techniques. In particular, concerning contrastive learning methodologies that heavily lean on large batch sizes, ablation experiments are conducted to ascertain the influence of batch size on our occluded image contrastive learning approach, as shown in Table 3. The utilization of the ViT-L/16 model for validation reveals that a larger effective batch size correlates with improved model performance and efficiency, aligning with the consensus within the community [19, 20]. We attribute this to the statistical advantage of expanded negative sample pools, which achieve broader coverage of the latent feature space. Such comprehensive sampling sharpens inter-class separation by refining decision boundaries during contrastive optimization. However, limited by computational resources, we are unaware of the maximum capacity of the effective batch size.
@@ 3.2.3 MLP Head is Not You Need
Table 4: Ablation experiments on MLP head. The results are based on the ImageNet-1K with the ViT-L/16. All methods are evaluated by pre-training hours, linear probing (LIN) and fine-tuning (FT). Correspondingly, we give different effective batch sizes related to the MLP head within the pre-trained model. Besides, pre-training hours on A100 are provided. The resolution of images is fixed to 224×224. Top-1 accuracy is used as the metric.
In many contrastive learning methods [1, 12], the ViT model is typically paired with an MLP head [21] to assist with the pretext task. The ViT learns semantic features from the image, while the MLP head handles the pretext classification task, enhancing the ViT’s learning capabilities, as shown in earlier work. However, traditional contrastive learning methods rely on handcrafted data augmentations to generate and process multiple image views at the instance level. In contrast, OCL explores an alternative approach: using masked images to create diverse views with fine-grained conceptual differences at the token level. As a result, OCL does not require an MLP head for instance-level classification tasks.
The ablation experiments on the MLP head involve testing different configurations, including a 2-layer MLP, a 3-layer MLP, and no MLP head (see Table 4). The design of the MLPs follows MoCo v3 [12], where the hidden layers of both the 2-layer and 3-layer MLPs are 1024-dimensional and employ the GELU activation function [22]. The output layers of both MLPs are 512-dimensional and do not use GELU. Additionally, all layers in the MLPs incorporate Batch Normalization (BN) [23], consistent with the methodology in SimCLR [11]. Due to GPU memory constraints, the batch size was adjusted to 1,800 for validation. From the fine-tuning results presented in Table 4, OCL is capable of operating effectively without an MLP head. Unlike traditional contrastive learning approaches, the absence of an MLP head does not degrade the model’s performance. On the other hand, the inclusion of additional auxiliary modules, such as deeper MLP heads, increases both the cost and time required for pre-training. However, the benefits of these modules do not outweigh the performance gains achieved by simply increasing the batch size.
@@ 3.2.4 Concentration of Similarity Metric $\kappa$
Table 5: Ablation experiments on concentrate parameter of kappa. The results are based on the ImageNet-1K with the ViT-L/16. Fine-tuning (FT) results are provided. The resolution of images is fixed to 224×224. Top-1 accuracy is used as the metric.
Following the approach of T-SP [14], we conduct ablation experiments to evaluate how different concentrations of the T-distributed adapter affect the pre-training model’s performance. The results are shown in Table 5. This study involves five specific degrees of concentration, namely $\kappa=4$, $16$, $32$, $64$, and $128$. From the Table 5, we can infer that as kappa increases, the pre-training performance of the model improves progressively until reaching $\kappa=64$ with the fine-tuning result of 85.8 under ViT-L/16. We explain that as kappa increases, the model must extract more precise semantic concepts from positive samples, creating a more challenging pretext task for pre-training. However, if kappa becomes too large, the model converges too slowly, which negatively impacts the overall effectiveness of the pre-training process.
@@ 3.3 Comparisons with previous results
BEiT w/o 800 - 83.2 - 85.2
MAE w/o 1,600 68.0 83.6 76.0 85.9
CAE w/o 1,600 70.4 83.9 78.1 86.3
I-JEPA w/o 600 72.9 - 77.5 -
Contrastive Learning
DINO w/ 1,600 78.2 82.8 - -
MoCo v3 w/ 600 76.7 83.2 77.6 84.1
@@
SiameseIM w/ 1,600 78.0 84.1 - -
ccMIM w/o 800 68.9 84.2 - -
ConMIM w/ 800 - 85.3 - 86.5
iBOT w/ 1,600 79.5 - 81.0 84.0
OCL w/o 1,600 74.2 83.4 77.9 85.8
@@
To demonstrate OCL’s ability to learn high-level conceptual representations without relying on handcrafted data augmentations, we compare its performance on linear probing and fine-tuning tasks under pre-training on ImageNet-1k. Importantly, while masked images are used to reduce semantic redundancy, OCL avoids reconstructing masked images. Instead, it uses contrastive loss to guide the network’s learning process. As a result, OCL is categorized as a contrastive learning method.
@@
Table.6 exhibits the performance of our method under the fine-tuning and linear probing for ImageNet-1k classification. Our OCL method has demonstrated outstanding performance in contrastive learning, notably obviating the need for intricate augmentations across multiple views. It validates the practicality and efficacy of masked images for generating diverse views encapsulating distinct fine-grained semantic concepts, which diminishes conceptual redundancy and expedites the conceptual pre-training process. Beyond that, it also stands out the significant competency of contrastive learning in extracting high-level semantic concepts.
@@
Compared to MIM methods, due to the absence of pixel-level image reconstruction in OCL, the training process is finished in 800 epochs, in contrast to the 1600 epochs required for MAE [2], CAE [24]. It corroborates the contribution of our methodology from the perspective of efficiency and high-level semantic concept extraction. Moreover, our OCL method is superior to BEiT [3] regarding the fine-tuning and linear probing results, also confirming the competency and robustness of our method.
@@
Concerning the fusion of MIM and CL, significant endeavours [25, 26, 27] are made to enhance pre-training performance on downstream tasks. However, they result in inefficiencies due to increased training iterations, augmented data, auxiliary modules, and diverse loss combinations. For instance, SiameseIM [25], ConMIM [26] and iBOT [27] leverage both hand-crafted data augmentation and more training epochs. Despite our model also integrating masking strategy with contrastive learning, it does not rely on hand-crafted view data augmentations and additional auxiliary modules, reconciling the divergence between efficient visual representation and effective conceptual pre-training. Besides, our model OCL also achieves competitive results on downstream tasks of linear probing and fine-tuning, and more importantly, shows promising scaling behavior. More comparison results of downstream tasks are provided in the supplementary.
Methods Epochs COCO ADE20k
$AP^{b}$ $AP^{m}$ mIoU
ViT-B/16
Supervised [25] 300 47.9 42.9 47.4
DINO [13] 800 50.1 43.4 46.8
iBOT [27] 1600 51.2 44.2 50.0
DenseCL [28] 400 46.6 41.6 44.5
MoCo-v3 [29] 600 47.9 42.7 47.3
BEiT [3] 800 49.8 44.4 47.1
MAE [2] 400 50.6 45.1 45.0
MAE [2] 1600 51.6 45.9 48.1
GreenMIM [30] 800 50.0 44.1 -
EsViT [31] 300 - - 47.3
MixMAE [32] 300 52.3 46.4 49.9
MixMAE [32] 600 52.7 47.0 51.1
SiameseIM [25] 400 50.7 44.9 49.6
SiameseIM [25] 1600 52.1 46.2 51.1
SimMIM [33] 300 51.1 45.4 48.9
OCL 800 51.5 45.5 46.1
ViT-L/16
MoCo-v3 [29]
BEiT [3] 800 53.3 47.1 53.3
MAE [2] 1600‘ 53.3 47.2 53.6
SimMIM [33] 800 53.8 - 53.6
MixMAE [32] 600 54.3 48.2 53.8
OCL 800 53.2 47.0 53.2
Table 7: Comparsion with previous methods on various downstream tasks, including object detection and segmentation on COCO and ADK20K. We report $\text{AP}^{\text{box}}$ ($\text{AP}^{\text{b}}$) and $\text{AP}^{\text{mask}}$ ($\text{AP}^{\text{m}}$) on COCO, and mIoU on ADE20K. Arch. represents the model architecture, where ViT-B/16 and ViT-L/16 are utilized to validate the performance of various methods.
Beyond that, we conducted supervised fine-tuning on the COCO dataset for object detection and instance segmentation using the Mask RCNN [34] framework, with our pre-trained encoder serving as the backbone. We follow the setup of MixMAE [32], with the window size to 16 $\times$ 16 to align with the 1024 $\times$ 1024 input image resolution. In terms of semantic segmentation on the ADE20k dataset, We use the UperNet [35] framework with our pre-trained encoder as its backbone, with the changed window size as mentioned above. The results are shown in Table 7, that our model achieves competitive results compared to other models with efficient pre-training progress. Compared to the contrastive learning method of MoCo-v3, our method obtains 3.6% improvement of $\text{AP}^{\text{box}}$ under ViT-B/16, especially without view data augmentations. Besides, the consistent performance gains on ADE20K semantic segmentation, achieving +7.1 mIoU improvements from ViT-B to ViT-L, empirically validate our method’s scalability across model capacities while maintaining parameter efficiency
Methods Epochs IN-A IN-R IN-S
MSN [36] 1200 37.5 50.0 36.3
iBOT [27] 1600 42.4 50.9 36.9
DenseCL [28] 400 30.8 43.8 29.9
MoCo-v3 [29] 600 32.4 49.8 35.9
MAE [2] 1600 35.9 48.3 34.5
SiameseIM [25] 1600 43.8 52.5 38.3
OCL 800 42.2 52.3 37.6
Table 8: Comparsion with previous methods on generalization capability and robustness on ImageNet-A (IN-A), ImageNet-R (IN-R) and ImageNet-Sketch (IN-S) datasets. ViT-B/16 is utilized as the backbone to validate the performance of various methods. Top-1 accuracy is used as the metric.
Beyond that, we conduct experiments in the domain generalization setting following the SiameseIM [25], as shown in Table 8. Experiments are conducted on ImageNet-A [37], ImageNet-R [38] and ImageNet-Sketch [39], with ImageNet [17] as the source dataset. Our method achieves results with a competitive edge, demonstrating the advantages of our model in generalization and robustness. Notably, our framework achieves competitive performance with significantly fewer training epochs than conventional contrastive approaches, demonstrating remarkable training efficiency while maintaining stable convergence—a critical advantage for scaling to large-scale datasets and complex architectures.
4 Conclusions and Outlooks
@@
In this paper, we introduce OCL, a novel, simple, and effective pre-training paradigm for visual conceptual representation. Our approach uses a masking strategy to generate diverse views with fine-grained semantic differences, enabling contrastive learning to classify and learn agreements within a mini-batch. This design eliminates the need for I) addressing semantic conceptual redundancy within images, II) reconstructing images, III) hand-crafted data augmentations, and IV) additional auxiliary modules, thereby improving efficiency and scalability. Experiments showcase the efficiency and scalability of our method, yielding competitive results compared to previous approaches. Additionally, ablation experiments provide insights that could inspire future pre-training paradigms.
We hope that our work will inspire future advancements in contrastive pre-training paradigms, specifically efficient visual representation. Our further work will focus on enhancing computational efficiency for billion-parameter models without sacrificing representation quality.
Appendix A Related Works
A.1 Masked Image Modeling
Inspired by masked language modeling in NLP, the core of masked image modeling is to predict the masked part of the input image. Among them, BEiT [3] tokenizes image patches through the reconstruction of the individual image using dVAE, and then predicts the tokens of the masked patches to learn visual representation. Similarly, MAE [2] utilizes a high masked ratio (75%) to corrupt the image and directly reconstruct the pixel-level masked image patches. Subsequently, numerous studies have referenced this paradigm for pre-training endeavours, including DropPos [5], U-MAE [8] and CAE [24]. DropPos [5] incorporates position reconstruction to bolster the spatial awareness of ViTs. U-MAE [8] introduces a uniformity loss as a regularization to the MAE loss to further encourage the feature consistency of the pre-training, and addresses the dimensional feature collapse. CAE [24] decouples the learning processes for image representation and pretext tasks, enabling the pre-trained model to prioritize image representation while disregarding the pretext task. Additionally, following the successful implementation of MAE, its applicability has been extended to diverse disciplines, such as SiamMAE [9] and MR-MAE [40]. Similarly, I-JEPA [18] predicts the feature encoding of the contextual region via MIM, and employs contrastive learning to align the features of neighbouring regions at the feature level.
Besides, latentMIM [41] tackles key training challenges in Latent MIM, showcasing its ability to produce spatially diverse, high-level semantic representations. Context Autoencoders [42] employ an encoder-decoder framework optimized with a combined reconstruction loss and alignment constraint, ensuring predictable representations of missing patches. data2vec [43] predicts representations of missing patches using an online target encoder, eliminating the need for handcrafted augmentations and achieving strong performance across vision, text, and speech modalities. Its successor, data2vec-v2 [44], further explores efficient architectures for multimodal learning.
A.2 Contrastive Learning
Among these methods, SimCLR [1] relies on complex pre-processing techniques to create distinct views of an image, aiming to make the task challenging enough to learn effective visual representations during pre-training. However, it requires large batch sizes to generate negative sample pairs, leading to long training times and high computational costs. DINO [13] uses a student-teacher architecture to extract visual representations from different views. Similarly, MoCo v3 [12] employs momentum updates to optimize its auxiliary network. These methods highlight that the core of contrastive learning lies in creating views with significant differences. However, this is challenging due to the inherent semantic redundancy in images. Additionally, ConCL [45] generates distinct concepts by cropping images and applies contrastive learning within a teacher-student framework, specifically for pre-training on pathological images.
DenseCL [28] introduces the concept of dense contrastive loss, which calculates the contrastive loss between dense feature vectors generated by the dense projection head at the local feature level, contrasting with traditional contrastive learning methods that operate at the global feature level. Moreover, MaskCo [46] utilizes the teacher-student network for generating image features of query and keys. The key features of different images are used to query certain images to generate positive and negative examples for contrastive learning. Furthermore, MSN [36] leverages data augmentation to create two views known as the anchor view and the target view. Subsequently, a random mask is applied to the anchor view, aligning the representation of the masked anchor view with the clusters of the unmasked target view.
A.3 Combination between MIM and CL
There are also a lot of efforts to bridge the gap between mask image modeling and contrastive learning, where the integration of teacher-student models emerges as the prevailing approach. iBOT [27] employs the teacher-student network to independently encode the two augmented views, with the student network processing masked images. The objectives of MIM and CL are jointly trained for self-distillation. Likewise, MST [47] introduces a masked token strategy leveraging multi-head self-attention maps, which selectively mask the tokens of the student network based on the output self-attention map of the teacher network, ensuring vital foreground remains intact. Similarly, SiameseIM [25] employs a Siamese network featuring two branches. The online branch encodes the initial view and predicts the representation of the second view based on their relative positions. Meanwhile, the target branch generates the target by encoding the second view. MSCN [48] generates multiple augmented views from input images and applies random masking. These masked views are encoded using a standard ConvNet, with representations optimized through a joint-embedding loss. RECON [49] integrates generative and contrastive learning paradigms via ensemble distillation, where a generative student model guides a contrastive student to unify both approaches. CMAE [50] employs a dual-branch architecture: an online branch with an asymmetric encoder-decoder for reconstructing masked images, and a momentum branch with a momentum encoder for contrastive learning on full images. This design enables holistic feature learning and enhanced discrimination.
ccMIM [51] leverages a contrastive loss to aid the reconstruction task as a regularizer, facilitating the extraction of image-wide global information from both masked and unmasked patches. Likewise, ConMIM [26] produces simple intra-image inter-patch contrastive constraints as the sole learning objectives for masked patch prediction, and strengthens the denoising mechanism with asymmetric designs to improve the network pre-training. Additionally, CoMAE [52] also applies CL to assist cross-modal MIM tasks. Besides, LGP [53] integrates MIM and CL in a sequential cascade manner: early layers are first trained under one MIM loss, on top of which latter layers continue to be trained under another CL loss.
A.4 Differences of OCL from Existing Methods
Masking has been adopted as an effective data augmentation technique to enhance training efficiency in several studies [54, 55, 56, 57, 58]. Previous studies have demonstrated its effectiveness in improving model performance and training efficiency [54, 55]. ExtreMA [57] employs random masking as a computationally efficient augmentation for Siamese representation learning, accelerating learning and enhancing performance on large datasets. MSN [58], the most relevant work to ours, generates two image views: a masked anchor view and an unmasked target view, aiming to cluster their representations. CAN [56] integrates contrastive learning, masked auto-encoding, and diffusion denoising into a unified framework. Our approach distinguishes itself from previous hybrid methods by achieving superior performance with a better performance-efficiency trade-off. Specifically, our masking strategy generates semantically diverse views and leverages contrastive learning to promote classification agreement within mini-batches. This approach eliminates: (I) semantic redundancy, (II) image reconstruction, (III) hand-crafted augmentations, and (IV) additional auxiliary modules, resulting in enhanced efficiency and scalability.
Appendix B Experiments
In this section, we first introduce the utilized datasets for pre-training and various downstream tasks. Subsequently, implementation details are provided. Finally, we present more results on downstream tasks, such as object detection, segmentation and domain generalization.
B.1 Datasets
The ImageNet-1k dataset [17] is a widely used image dataset consisting of 1.28M labeled images across 1k categories, with 50K validation images and 100k test images. The dataset has been instrumental in advancing computer vision research by providing a large-scale benchmark for image classification tasks. By leveraging the vast amount of labeled images in ImageNet-1k, self-supervised models can learn rich representations of visual data in an unsupervised manner, which can then be fine-tuned on downstream tasks with smaller labeled datasets. Besides, ImageNet-A [37], ImageNet-R [38] and ImageNet-Sketch [39] are leveraged for validation of the generalization capability and robustness of the vision model, with the training source of ImageNet.
MSCOCO [59] dataset is a large-scale dataset widely used for object detection and instance segmentation tasks created by Microsoft Research Asia. The COCO Detection dataset contains more than 330K images, including more than 1.5M labeled object instances for 80 different categories. Each object instance is labeled with category, bounding box, and segmentation mask information.
The ADE20K [60] semantic segmentation dataset comprises over 20K scene-centric images for 150 semantic categories, meticulously annotated with pixel-level object and object parts labels.
B.2 Implementation Details
We employ ViT-Base and ViT-Large as our visual backbones, respectively. Among them, Vit-Base consists of 12 transformer encoder layers and an FFN intermediate size of 3,072. The hidden dimensions of the ViT-Base are 768, with 12 attention heads. The number of parameters is about 86 million. The input image size is set to $224\times 224$. In terms of ViT-L/16, ViT-L/16 consists of 24 transformer encoder layers and an FFN intermediate size of 4,096. The input image size is set to $224\times 224$, with a patch size of $16\times 16$. The hidden dimensions of the ViT-Large are 1,024, with 16 attention heads. And, the number of parameters is about 307 million.
In terms of the pre-training progress, the hyperparameters are presented in Table 9. We utilize the AdamW optimizer, which is configured with a cosine annealing schedule as the learning policy. The initial base learning rate is set to $1.5\times 10^{-4}$, and the AdamW optimizer is employed with hyperparameters $\beta=(0.9,0.95)$. Additionally, we set the weight decay to 0.05 without dropout. We use the strategy of cosine learning rate decay, with 40 warm-up epochs. Unless otherwise specified, the pre-training of our vision language model consists of 800 epochs, executed on $2\times 2$ NVIDIA A100 GPUs.
Table 9: The pre-training hyperparameters.
ViT-B/16 ViT-L/16
Training Epochs 800 800
Warmup Epochs 40 40
Optimizer AdamW AdamW
Base Learning Rate 1.5e-4 1.5e-4
Learning Rate Decay Cosine Cosine
Adam $\beta$ (0.9, 0.95) (0.9, 0.95)
Weight Decay 0.05 0.05
Eff. Batch Size 9,600 2,400
Concerning the downstream tasks of fine-tuning and linear probing on ImageNet, the hyperparameters are shown in Table 10 and Table 11.
Table 10: The fine-tuning hyperparameters.
ViT-B/16 ViT-L/16
Training Epochs 100 50
Warmup Epochs 5 5
Optimizer AdamW AdamW
Base Learning Rate 5e-4 1e-3
Learning Rate Decay Cosine Cosine
Adam $\beta$ (0.9, 0.95) (0.9, 0.95)
Weight Decay 0.05 0.05
Eff. Batch Size 1,024 1,024
Table 11: The linear probing hyperparameters.
ViT-B/16 ViT-L/16
Training Epochs 90 50
Warmup Epochs 10 10
Optimizer LARS LARS
Base Learning Rate 0.1 0.1
Learning Rate Decay Cosine Cosine
Weight Decay 0.0 0.0
Eff. Batch Size 16,384 1,024